OpenStudy (anonymous):

give the slope for 5x+9y=0

OpenStudy (anonymous):

solve for y, and the slope is the number next to x

OpenStudy (anonymous):

5x+9y=0 -5x -5x 9y=5x divided both sides by 9 y = (5/9)x. The slope is (5/9)